Multiply 62 by 35, firstly by breaking down 35 in its terms (30 + 5) and secondly by breaking down 35 in its factors (5 x 7). Show all your steps.

(a) 62 x (30 + 5)

(b) 62 x (5 x 7)


1
Expert's answer
2022-06-06T13:30:51-0400

a)


"62\times 35=62\times(30+5)=62\times 30+62\times5"

"=1860+310=2170"

b)


"62\times 35=62\times(5\times 7)=(62\times 5)\times7"

"=310\times 7=2170"